OpenStudy (anonymous):

find the exact value of sec(pi/8)

OpenStudy (freckles):

\[\frac{1}{2} \cdot \frac{\pi}{4}=\frac{\pi}{8}\]

OpenStudy (freckles):

1/2*pi/4 means half of pi/4 think half angle identity

OpenStudy (anonymous):

I'm confused where the half angle identity comes in

OpenStudy (freckles):

well we know pi/4 is on the unit circle and half of pi/4 is pi/8 so since pi/8 is in the first quadrant we know cos to be positive \[\cos(\frac{\pi}{8}) \\ =\cos(\frac{1}{2} \cdot \frac{\pi}{4})=\sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}}\]

OpenStudy (anonymous):

I haven't seen that equation before

OpenStudy (freckles):

\[\cos(\frac{\theta}{2})= \pm \sqrt{\frac{1+\cos(\theta)}{2}}\]? you haven't seen this one?

OpenStudy (freckles):

have you seen this one: \[\cos^2(\theta)=\frac{1}{2}(1+\cos(2 \theta))\]

OpenStudy (anonymous):

maybe I wrote it down for the wrong thing

OpenStudy (anonymous):

My notes are all wrong lol

OpenStudy (freckles):

http://www.purplemath.com/modules/idents.htm here are a lot of trig identities with the corresponding names

OpenStudy (freckles):

maybe you can use this to fix your notes

OpenStudy (anonymous):

\[\cos \frac{ x }{ 2 }=\sqrt{\frac{ 1+sinx }{ 2 }}\] that's what I have

OpenStudy (anonymous):

+- in the front

OpenStudy (freckles):

err yes that sin(x) thing is definitely suppose to be cos(x)

OpenStudy (anonymous):

Apex is terrible

OpenStudy (freckles):

\[\cos(\frac{x}{2})= \pm \sqrt{\frac{1+\cos(x)}{2}} \\ x=\frac{\pi}{4} \\ \cos(\frac{\pi/4}{2})=\pm \sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}} \\ \\ \cos(\frac{\pi}{8}) \text{ is positive since } \frac{\pi}{8} \text{ is \in the first quadrant } \\ \cos(\frac{\pi}{8})= \sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}} \\ \sec(\frac{\pi}{8})=\frac{1}{\cos(\frac{\pi}{8})} =\sqrt{\frac{2}{1+\cos(\frac{\pi}{4})}} \]

OpenStudy (freckles):

there is still a couple things for you to do

OpenStudy (freckles):

find cos(pi/4) using unit circle

OpenStudy (freckles):

and simplify

OpenStudy (anonymous):

how'd you get cos pi/4 ?

OpenStudy (freckles):

x is pi/4

OpenStudy (freckles):

i replace x with pi/4

OpenStudy (freckles):

\[\frac{1}{2} \cdot \frac{\pi}{4} =\frac{\pi}{8} \\ \text{ so } x=\frac{\pi}{4}\]

OpenStudy (anonymous):

oooh okay

OpenStudy (anonymous):

cos pi/4 is sqrt2/2

OpenStudy (anonymous):

and 45 degrees

OpenStudy (freckles):

right so you can replace cos(pi/4) is sqrt(2)/2

OpenStudy (freckles):

well pi/4 is 45 deg

OpenStudy (freckles):

\[\cos(\frac{x}{2})= \pm \sqrt{\frac{1+\cos(x)}{2}} \\ x=\frac{\pi}{4} \\ \cos(\frac{\pi/4}{2})=\pm \sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}} \\ \\ \cos(\frac{\pi}{8}) \text{ is positive since } \frac{\pi}{8} \text{ is \in the first quadrant } \\ \cos(\frac{\pi}{8})= \sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}} \\ \sec(\frac{\pi}{8})=\frac{1}{\cos(\frac{\pi}{8})} =\sqrt{\frac{2}{1+\cos(\frac{\pi}{4})}} \\ \sec(\frac{\pi}{8})=\sqrt{\frac{2}{1+\frac{\sqrt{2}}{2}}}\]

OpenStudy (anonymous):

yeah I didn't know if I needed the degrees or radians so I put both lol

OpenStudy (freckles):

you don't need either you just need to find cos(45) or cos(pi/4)

OpenStudy (freckles):

cos(45 deg) is equivalent to finding cos(pi/4)

OpenStudy (anonymous):

do I need to get rid of the radical or leave that there ?

OpenStudy (freckles):

I would get rid of the compound fraction action inside the radical and then see if there are any perfect squares afterwards

OpenStudy (anonymous):

so would I multiply top and bottom by the opposite of the denominator ?

OpenStudy (freckles):

\[\sec(\frac{\pi}{8})=\sqrt{\frac{2}{1+\frac{\sqrt{2}}{2}}} \cdot \sqrt{\frac{2}{2}} \\ \sec(\frac{\pi}{8}) = \sqrt{ \frac{2 (2)}{(1+\frac{\sqrt{2}}{2})(2)}} \\ \sec(\frac{\pi}{8})=\sqrt{\frac{4}{2+\sqrt{2}}}\] notice the numerator inside the squareroot is a perfect square

OpenStudy (freckles):

\[\sec(\frac{\pi}{8})=\frac{\sqrt{4}}{\sqrt{2+\sqrt{2}}}\]

OpenStudy (anonymous):

wouldn't the rad 2 or rad 2 equal 1 ?

OpenStudy (freckles):

yes sqrt(2/2)=1 that is why I multiplied it to get rid of the compound fraction action inside the square root

OpenStudy (anonymous):

I'm confused on the multiplying of 2 in the denominator and numerator

OpenStudy (freckles):

let's look inside the square root we have inside the square root: \[\frac{2}{1+\frac{\sqrt{2}}{2}}\] this has a mini-fraction the sqrt(2)/2 the sqrt(2)/2 has a 2 in the denominator I multiply top and bottom inside the square root by 2 so I have inside the square root: \[\frac{2}{1+\frac{\sqrt{2}}{2}} \cdot \frac{2}{2} \\ \frac{2(2)}{(1+\frac{\sqrt{2}}{2})(2)} \\ \frac{4}{1(2)+\frac{\sqrt{2}}{2}(2)} \\ \frac{4}{2+ \frac{2}{2} \sqrt{2}} \\ \\ \frac{4}{2+(1) \sqrt{2}} \\ \frac{4}{2+ \sqrt{2}}\]

OpenStudy (anonymous):

why did you transfer the radical to the two ?

OpenStudy (freckles):

what 2 ?

OpenStudy (freckles):

are you asking me about this: \[\frac{\sqrt{2}}{2}(2)=\frac{2}{2} \sqrt{2}\]

OpenStudy (anonymous):

this part

OpenStudy (freckles):

multiplication is commuative

OpenStudy (freckles):

\[\frac{\sqrt{2}}{2}(2) =\frac{\sqrt{2}}{2} \cdot \frac{2}{1} = \frac{\sqrt{2} \cdot 2 }{2 \cdot 1 } \\ \text{ you change order \in multiplication } \\ =\frac{2 \sqrt{2}}{2 \cdot 1} \\ \text{ you can reseperate the fraction } \\ =\frac{2}{2} \frac{\sqrt{2}}{1} \\ =1 \frac{\sqrt{2}}{1} \\ 1 \cdot \sqrt{2} \ = \sqrt{2}\]

OpenStudy (anonymous):

oooh okay , that makes sense

OpenStudy (freckles):

all I did was use the fact that a*b is the same as b*a

OpenStudy (anonymous):

okay, that makes a lot of sense

OpenStudy (anonymous):

for my final answer , I got \[-\sqrt{2}+2\] is that right ?

OpenStudy (freckles):

\[\cos(\frac{x}{2})= \pm \sqrt{\frac{1+\cos(x)}{2}} \\ x=\frac{\pi}{4} \\ \cos(\frac{\pi/4}{2})=\pm \sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}} \\ \\ \cos(\frac{\pi}{8}) \text{ is positive since } \frac{\pi}{8} \text{ is \in the first quadrant } \\ \cos(\frac{\pi}{8})= \sqrt{\frac{1+\cos(\frac{\pi}{4})}{2}} \\ \sec(\frac{\pi}{8})=\frac{1}{\cos(\frac{\pi}{8})} =\sqrt{\frac{2}{1+\cos(\frac{\pi}{4})}} \\ \sec(\frac{\pi}{8})=\sqrt{\frac{2}{1+\frac{\sqrt{2}}{2}}} \\ \] \[\sec(\frac{\pi}{8})=\sqrt{\frac{2}{1+\frac{\sqrt{2}}{2}}} \cdot \sqrt{\frac{2}{2}} \\ \sec(\frac{\pi}{8}) = \sqrt{ \frac{2 (2)}{(1+\frac{\sqrt{2}}{2})(2)}} \\ \sec(\frac{\pi}{8})=\sqrt{\frac{4}{2+\sqrt{2}}}\] \[\sec(\frac{\pi}{8})=\frac{\sqrt{4}}{\sqrt{2+\sqrt{2}}}\] \[\sec(\frac{\pi}{8})=\frac{2}{\sqrt{2+\sqrt{2}}}\]

OpenStudy (anonymous):

wouldn't the rad 2 cancel out since it's two radicals ?

OpenStudy (anonymous):

so it'd be \[\frac{ 2 }{ \sqrt{2}+2 }\]

OpenStudy (freckles):

no \[\sqrt{a+\sqrt{b}} \neq \sqrt{a}+b \\ \text{ take } a=4 \text{ and } b=16 \\ \text{ so } \sqrt{a}=2 \text{ and } \sqrt{b}=4 \\ \sqrt{a+\sqrt{b}}=\sqrt{4+4}=\sqrt{8} \\ \text{ while } \sqrt{a}+b=2+16=18 \\ \text{ and } 18 \text{ is definitely not equal to } \sqrt{8} \\ \text{ so } \sqrt{a+\sqrt{b}} \text{ is not } \sqrt{a}+b\]

OpenStudy (anonymous):

so it stays like that ?

OpenStudy (freckles):

it stays like \[\frac{2}{\sqrt{2+\sqrt{2}}}\] yes you could rationalize the bottom but it will still look ugly

OpenStudy (freckles):

so I would leave it as is

OpenStudy (anonymous):

I never learned that lol . I have to get ready for work now though , thank you for the help !

OpenStudy (freckles):

good luck

OpenStudy (anonymous):

If I post up a new question with this activity I have to do , do you think you could figure it out ? I tried but I couldn't get it . I got stuck

OpenStudy (anonymous):

The activities that are worth points in this packet are super confusing